OpenStudy (anonymous):

which of the following points are not on the graph on y=-2x+3?

OpenStudy (ganpat):

following points ??

OpenStudy (anonymous):

the answer choices are (1,1),(-2,3),(-1,5),(2,-1)

OpenStudy (ganpat):

so what u can is, put the points in equations.. if it satisfies.. than thats ur answer.. Inside the bracket, they are arranged as (x , y).. start doing for each one..

OpenStudy (anonymous):

they dont have a graph! how do u figure it out in ur head

OpenStudy (ganpat):

u dont need graph for this.. i will show u an example.. given equation = y=-2x+3 , and consider point (2, -1) So you have x = 2 and y = -1.. substitute in equation.. so -1 = -2 (2) + 3 -1 = -4 + 3 -1 = -1 as LHS = RHS... so this point is on the graph... similarly solve for others and the one in which LHS not equal to RHS is ur answer...

OpenStudy (anonymous):

so we figure it out it would be a 1,1 do u think its ritee?

OpenStudy (ganpat):

nopes.. it fits in the equation... 1 = -2 (1) + 3 1 = 1 the answer is 2, (-2,3) 3 = -2(-2) + 3 3 = 7.. so it will not fit in the equations and is not on the graph..

OpenStudy (anonymous):

can u explaion it step by step on how u got the answer?

OpenStudy (ganpat):

whats the equation we have ?? y=-2x+3.. and what is the point we have to check ?? (-2, 3)

OpenStudy (ganpat):

now compare this point with, (x, y ) and substitute in equation to solve further..